IOO meeting 2019 1007
Ohakawa, Sugai, Uehara, Takahashi, Fujikawa, Tanaka, (Akutsu), Somiya,
* MZM paper was approved. *
- The beam position on the MCo is off-centered by 1.5 cm or more.
- We might leave it by O3. It needs to vent the chamber to fix it.
- The input beam seems to be clipped at the IFI. We are working on fixing it.
OMC, OMMT, OSTM
- OMC commissioning is ongoing. (somiya)
At Sep. 25th. The jig was manufactured and tested. However, the PZT cable cannot be connected with the jig, and it is modified now. ->
The QPD driver did not work well, however, the driver itself seemed fine. -> Test with the LED.
- PZT driver is not delivered yet.
- OMC glue jig?
- PD mount will be delivered ~10/18.
- Curved steering mirror will be delivered ~end Oct.
- OSTM replacement (takahashi)
- Next week, the suspension will be assembled and tested with the dammy mirror.
- The mirror is delivered to NAOJ, and Matteo is working in the inspection.
- Installation schedule arrangement:
- OSTM should be installed after the OMC.
- Beam dump?
- RF PD has small power and it might limit the sensitivity soon.
- Double demodulation
- Scattering measurement
- Have another meeting on 10/8. Detail will be discussed there.
- June Gyu Park from Oct. 4th to
- Anamaria from LIGO ?? in Nov.?
- Nakano to LLO ~2 weeks in Oct.
- 11/11 (Mon.)
- 12/2 (Mon.)